Which equation is true for the function h(x) = 16^x?
h(–1.25) = –0.5
h(–0.5) = 0.25
h(0.75) = 12
h(1.25) = 20

Answer by stanbon(75887)

Which equation is true for the function h(x) = 16^x?
h(–1.25) = –0.5
16 to ANY power cannot be negative
-------------------------------------------
h(–0.5) = 16^(-1/2) = 1/sqrt(16) = 1/4 = 0.25::: True
---------------------------------
h(0.75) = 16^(3/4) = 2^3 = 8 # 12::: False
-----------------------------------
h(1.25) = 16^(5/4) = 2^5 # 20::: False
============
Cheers,
Stan H.
------------